OpenStudy (dls):

Example of any 2 particles which do not exert Gravitational force on each other?

OpenStudy (dls):

@experimentX @Mashy

OpenStudy (anonymous):

photons? :P

OpenStudy (dls):

I thought of photon but its not exactly massless..it has SOME mass so negligible..but no 0

OpenStudy (dls):

timing:D

OpenStudy (anonymous):

ermm.. is it?.. i thought it had no mass :D

OpenStudy (dls):

\[\LARGE M_{p}=1.8 \times 10^{-42}g\]

OpenStudy (anonymous):

omg.. :O.. where did you get that from? :O.. anwyays.. there are no particles then without gravity as far as i know :O

OpenStudy (dls):

if photons dont have mass how do they have momentum?

OpenStudy (anonymous):

that is really more subtle than you think.. when light can act like both particle and wave... which itself is mind boggling.. you think it not having mass but having momentum is a big deal? :P

OpenStudy (dls):

lol

OpenStudy (anonymous):

You measure gravitational force on particles that's at a quantum level, which involves photons. In QM photons are considered to have a rest mass of 0, so I guess the only answer is photon. Higgs to come?

OpenStudy (dls):

Higgs is massless?

OpenStudy (anonymous):

DLS look into photoelectric effect derivation and compton scattering

OpenStudy (anonymous):

but you know what?? think of this .. if two photons of light are moving parallel to each other(in vacuum)... if they put force of gravity on each other, then their velocities WOULD CHANGE right.. i don't care how small that change is.. but it would change.. ?? but remember relativity?? light cannot have ANY OTHER speed than C.. in vacuum.. regardless of what crap you do!!.. so they cannot put gravitational force on each other :O hmm.. how is that argument? :O

OpenStudy (anonymous):

for why it has momentum, no higgs would give more knowledge, it gives mass.

OpenStudy (dls):

i know photoelectric effect

OpenStudy (anonymous):

or maybe not.. you can use general relativity to explain that.. so never mind :D

OpenStudy (dls):

The Standard Model's explanation of why some fundamental particles have mass when 'naive' theory says Higgs boson should be massless

OpenStudy (dls):

so can we say photon photon and boson boson :O

OpenStudy (dls):

@Mashy dude stop showing off relativity in every question of mine when u know i dont understand it :P

OpenStudy (anonymous):

Photoelectric determines relationship between energy and frequency ... E proportional to (a + bw) where w = angular omentum so E = Hbar(w-w_a) = h(v - v_a) photon rest mass is zero for general relation of energy, so it can be written as E^2 = (m_0*c^2)^2 + p^2c^2 = Hbar^2w^2 and can cancel to momentum p = Hbar * k, where k = w/c (wave number) therefore the momentum of a photo can be given as p = Hbar * k bear in mind wave particle duality, how a wave can have momentum is a toughy, yet quite beautiful

OpenStudy (anonymous):

wait for @Iamgmg90 something bombastic is on its way.. either that or his computer is stuck :D

OpenStudy (anonymous):

ok sorry DLS :P.. didn't mean to!

OpenStudy (anonymous):

Bombastic haha, its just my understanding I hope it shed light on the matter, excuse the pun :P

OpenStudy (dls):

guys im a sweet little cute innocent student :3

OpenStudy (dls):

simple answers should be awesome :P

OpenStudy (anonymous):

Trying to word equations in english and failing, I don't think quantum mechanics could be simple hehe, are you a student of physics?

OpenStudy (dls):

yes..XIth grade

OpenStudy (anonymous):

11th grade? I have no idea what that is, I'm English

OpenStudy (anonymous):

is that high school?

OpenStudy (dls):

kinda

OpenStudy (anonymous):

I think the main part of understand that is knowing planck's constant well

OpenStudy (anonymous):

& the difference between h and hbar

OpenStudy (dls):

yes,ive read about atomic struc

OpenStudy (dls):

not hbar

OpenStudy (anonymous):

its just h / 2pi

OpenStudy (anonymous):

but not just, very important for its relationship with angular frequency

OpenStudy (anonymous):

omg sorry in those equations w = angular frequency, not momentum

OpenStudy (dls):

\[mvr=\frac{nh}{2\pi}\]

OpenStudy (dls):

:o

OpenStudy (anonymous):

Bohr is far from QM, but orbital win

OpenStudy (dls):

anyway,thanks guys! ill go with photon for meanwhile :P and/or bosson

OpenStudy (anonymous):

from where Bosson came?? bossons have mass :O !!!

OpenStudy (dls):

http://prntscr.com/tqp5w

OpenStudy (dls):

naive theory \m/

OpenStudy (anonymous):

so? it still says higg's particle gives mass!

OpenStudy (dls):

okay photon :P

OpenStudy (anonymous):

higg's particle has been discovered hello!??

OpenStudy (dls):

oops I cant give medal to myself D:

OpenStudy (anonymous):

super large hardron collider??

OpenStudy (dls):

oout off my syllabus :P

OpenStudy (anonymous):

geneva switzerland?? .. 7.7TEV??? hello?!?!

OpenStudy (dls):

:| thanks everyone